# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ III-B1 Coherent vs incoherent ONN
Here, $\kappa$ and $t$ represent the field coupling factor and transmission factor, respectively. $\alpha_{rt}=exp\left(-\alpha\cdot 2\pi R\right)$ and $\phi_{rt}=\beta\cdot 2\pi R$ are the round-trip field attenuation factor and phase, with $\alpha$ and $\beta$ being the real and imaginary parts of the complex transmission coefficients [107]. The resonance condition is $\phi_{rt}=2m\pi$, i.e., $\lambda_{\text{res}}=2\pi n_{\text{eff}}R/m$, where $m$ is a positive integer. A potential issue arises in achieving full-range modulation, ranging from 0 to 1, can only be attained with symmetric, lossless coupling ($\alpha_{rt}=1$, $t_{1}=t_{2}$), which is unfeasible for actual devices. An all-pass MRR can be regarded as a special case of add-drop MRR, distinguished by the absence of a drop bus waveguide with $t_{2}=1$ and $\kappa_{2}=0$ for Eq. (15). While the phase at through/drop port can also be derived from Eq. (15) and Eq. (16), the phase response and tuning range are highly sensitive to coupling conditions and round-trip losses. Hence, MRRs are primarily employed for amplitude modulation within incoherent systems. Another challenge comes from the experimental perspective: the response of the MRR is susceptible to environmental factors such as temperature variations and vibrations, thereby impairing resolution. For a more in-depth theoretical analysis of MRRs, readers may refer to relevant literature and books [107, 108].
In tensor operations, MRR can modulate the transmission of each port by tunning coupling coefficients or the round-trip phase [109]. Unlike MZIs, MRRs demonstrate significant wavelength selectivity. An MRR with a moderate quality factor can achieve a resonance peak with a 3 dB bandwidth of a few hundred picometers. The narrow bandwidth makes MRRs particularly effective as tunable filters in WDM-based systems, where they are often arranged in series for the selective modulation of signals with different wavelengths. This application is exemplified by the concept of “weight banks” proposed by Tait et al. [110, 111], as shown in Fig. 6.d. In this architecture, photodetectors can spontaneously perform the dot product of operands encoded on different wavelengths. Another incoherent architecture is the crossbar array, which implements MVMs using its programmable transfer matrix of tunable switch array [112, 113]. Ohno, et al. demonstrated a 4$\times$4 add-drop MRR crossbar array for MVMs, with the corresponding schematic illustrated in Fig. 6.e [112]. Specifically, each wavelength coming from the row bus waveguide can be weighted and subsequently directed into a specific column by the MRR array. The weighted elements can then be collected and aggregated by a photodetector at the end of the column, thus implementing MVMs.
@@
Figure 6: Various implementations of photonic-electronic tensor cores. (a) Schematic of a 2$\times$2 MZI where the beam splitters can be implemented by multi-mode interferometer (MMIs) or directional couplers. An alternative configuration of MZI features a single input and output, with the splitter designed as a Y-branch. (b) An MZI-based coherent photonics tensor core for ONNs [29]. The weight matrix can be represented by SVD in the form $\bm{U\Sigma V^{\dagger}}$. Due to chip size and complexity considerations, this work only implements $\bm{U}$ and $\bm{\Sigma}$ on a single pass through the circuit. (c) Schematic and power spectral responses of an add-drop MRR with power coupling coefficients ${\kappa_{1}}^{2}={\kappa_{2}}^{2}=0.2$ (assuming the coupling junctions are lossless, i.e., $\kappa^{2}+t^{2}=1$) and 5% round-trip power loss. (d) The weight bank architecture utilizes MRRs as tunable filters for parallel modulation of WDM signals [111]. The series-connected MRRs share a single bus waveguide as through port, while the other waveguide can be shared or independent, enabling the selective addition or removal of wavelengths. (e) Schematic of 4$\times$4 MRR crossbar array demonstrated in Ref [112]. Input signals are first modulated via MZMs, and then distributed and weighted through the crossbar array, followed by a photodetector array that sums the weighted WDM elements. (f) The Schematic of an optical tensor core supports dynamic full-range general matrix multiplication based on WDM and coherent interference [114].
@@ III-B2 Static and dynamic weight encoding
The above-mentioned architectures topically map one operand of MVMs (typically a weight matrix W) onto hardware, executing multiplication through the transmission matrix of PIC. Therefore, the operational speed of an ONN is, in theory, only determined by the modulation rate of input signals x. However, practical implementations face several limitations. Primarily, deploying MVMs with large dimensions is challenging due to the cost, complexity, loss, and other scalability issues of PICs. While multiplexing techniques can enhance the parallelism of ONNs, time-domain hardware reuse is necessary to fulfill the computational requirements of intricate machine-learning tasks. This indicates that the “weight-static” architecture constitutes a bottleneck for large networks because of the huge gap between ultra-fast optical computing and slow mapping/reprogramming. To fully unleash the potential of optical analog computing, the critical role of dynamic encoding for specific tasks should be recognized. For dynamically-operated ONNs, two essential requirements must be satisfied. First, the parameters in the network need to be high-speed reprogrammable for efficient hardware reuse. This requires the modulators that operate at the gigahertz-level rate, exemplified by those based on field-effect tuning mechanisms. Secondly, dynamic operation mandates that parameter mapping and output reading be conducted “directly” without additional signal preprocessing or data correction. For instance, to map a 12×12 matrix on MZI-mesh framework, it requires $\sim$1.5 ms to perform SVD and phase decomposition on a CPU [114], which precludes dynamic operation due to the delay.
The importance of dynamic encoding is further highlighted by its compatibility with Transformer [7]. Transformers and unique attention mechanisms have attracted significant research interest in recent years due to their exceptional performance in natural language processing (NLP), machine vision (CV), and large-scale language models (LLM). Unlike weight-static architectures, Transformer employs the multi-head self-attention (MHA) mechanism within its encoder and decoder blocks, necessitating matrix multiplication involving two dynamic, full-range operands. Zhu et al. presented a dynamically-operated PTC for the first photonic Transformer accelerator, leveraging coherent light interference and WDM [114]. As shown in Fig. 6.f, the elements within the input vectors x and y are encoded at distinct wavelengths, and then add a $-90^{\circ}$ phase shift to one vector. Subsequently, through a directional coupler, two orthogonal signals can be recombined into the form (x $\pm$ y), facilitating the computation of the dot product via a balanced photodetector.
@@ III-B3 Hardware-efficient PTCs
For the aforementioned structures, the total number of optical components required to implement an $m\times n$ general matrix is similar, e.g., $m(m-1)/2+n(n-1)/2+\max(m,n)$ MZIs for SVD-based MZI mesh, and $m\times n$ MRRs in microring-based ONNs. In addition to the high hardware costs and the inherent control complexities, the required electrical components take a large proportion of energy consumption, particularly for high-speed and high-resolution modulation. To these challenges, researchers are exploring strategies to enhance hardware efficiency at multiple levels—from the device to the architecture—to improve the scalability of ONNs.
Firstly, the optical components or structures featuring compact topology have been proposed to reduce PIC footprint and improve the hardware efficiency for tensor operations. Zhu et al. developed an ONN architecture consisting of two diffractive cells and MZIs exhibiting linear scaling in relation to the input dimensions [115], as shown in Fig. 7.a. The star-shaped diffractive cells shows the capability of performing on-chip parallel Fourier transform and convolution operations. In a similar vein, an ONN leveraging on the combination of WDM technology and MMI has been developed and demonstrates an accuracy of 92.17% on the MNIST dataset. Beyond waveguide-based PIC, Wang et al. proposed a metasurface-based processing unit that provides ultra-high throughput for MVM [116] (Fig. 7.b). Within the subwavelength structure, each pair of slots acts as a weight element and connects to the following layers via in-plane diffraction and interference. For a specific neural network, the width and length of the slot need to be designed to map corresponding weights.
The task-specific passive devices not only yield a smaller footprint and enhanced hardware efficiency but also lower power consumption. However, a major challenge associated with passive device-based tensor cores is their fixed configuration, leading to potential effectiveness degradation across varying tasks. To address this challenge, active multi-operand devices present another opportunity to break the fundamental limitation to achieve high-density tensor operation by squeezing MAC operation into a single device. Specifically, the execution of the length-k dot product between input vector $\bm{x}_{\text{in}}$ and weight vector $\bm{w}$ by k-operand devices can be represented as follows:
@@
where the function $f(\cdot)$ represents the E-O transmission function, while $g(w,x)$ is related to the encoding mechanism. Here, $\bm{w}$ can be encoded using programmable resistances (such as memristors or PCMs), tunable amplifiers/attenuators, or the effective modulation length in cases, while $\bm{x}_{\text{in}}$ can be encoded by electrical signals. It is worth noting that, unlike conventional ONNs, the nonlinear transfer function of the device has to be considered during the training of models. The type of multi-operand devices can either be the modulators mentioned above or be based on other passive optical components such as MMI [117]. Feng et al. first demonstrated this method experimentally on a 4-operand MZI, achieving a measured accuracy of 85.9% in SVHN recognition tasks with 4-bit control precision (Fig. 7.c). Besides, Gu et al. proposed a compact ONN architecture based on multi-operand MRRs (MOMRR) [118]. The MOMRR-based ONN supports weight encoding through two sets of rails, with full-range results carried out by balanced photodetectors [119]. Additionally, the architecture can be combined with the structured pruning strategy to further to further improve network scalability. Theoretically, multi-operand optical synapses could execute vector operations with nearly the same footprint as the single-operand counterparts. However, the intrinsic nonlinearity and possible crosstalk among operands could present challenges for training and calibration.
@@
Beyond improvements at the device level, another promising approach enhances hardware efficiency at the circuit level by software-hardware co-design approaches. Subspace networks, for instance, sacrifice a portion of matrix representability in exchange for fewer parameters instead of implementing universal linear operations or general matrix multiplication (GEMM). The effectiveness of this strategy can extend to ONNs by trading the universality of weight representation for higher hardware efficiency. An example is butterfly-style photonics tensor core [81] (Fig. 7.d). By parameter pruning, this architecture could be implemented using significantly fewer MZI by pruning with a scale factor of $O(nlog_{2}^{n})$ rather than $O(n^{2})$ required by MZI mesh for GEMM. To evaluate the effectiveness of a machine learning task, this study experimentally showcases a measured accuracy of over 94% on the MNIST hand-written digits classification task, with up to 7$\times$ fewer active optical components, 3.3× smaller footprint, and 5.5$\times$ lower latency compared to conventional MZI mesh. For subspace networks, the trade-off between hardware efficiency and matrix representability is an important consideration, and more details will be discussed in Section V-A.
@@
Beyond the one-shot broadcast architecture, an alternative method leverages time-domain integration with fewer optical components to perform MACs [120, 121]. Following this approach, Sludds et al. developed a delocalized optical computing system that showcases edge computing capabilities over a span of 86 kilometers [120]. As shown in Fig. 7.e, the modulated WDM signals are separated by a passive demultiplexer and then fed to a set of time-integrating receivers. On the client side, only a single MZM, ADC, and DAC are used, which allows an ultra-low power assumption of fJ per MAC level. The photocurrent $I(t)$ generated by the photodetector produces a voltage across the integrating capacitor $C$ by accumulating charge, thereby achieving the summation operation, which can be expressed as Eq. (18).
@@
This strategy compromises the speed benefit inherent in optical computing to achieve a smaller chip footprint and reduced hardware complexity, which allows milliwatt-class edge devices.
@@ III-C Implementations of Nonlinearities
Activation functions introduce the necessary nonlinearity into the network, allowing it to learn and adapt to a wide range of complex tasks. For PIC-based photonic neurons, the implementations of nonlinearities fall into two major categories: the optical-electrical-optical (O-E-O) and the all-optical approach.
@@ III-C1 O-E-O nonlinearities
The activation function within the O-E-O neurons can be realized most directly by routing the weighted sum through an A/D conversion process into digital processing units, such as CPUs, GPUs, or FPGAs. This process involves performing nonlinear operations digitally and then converting the output back into analog signals to feed into photonic neurons. The main advantage of this approach is the extensive transfer ability of existing ANN architectures to the PIC platform, and it allows the implementation of arbitrary active functions. However, the compulsory A/D and D/A conversion, along with digital processing, impose significant latency and power consumption, thereby undermining the efficiency advantages inherent in ONNs. As mentioned in Section II, the performance and speed of an optical computing system are constrained by its weakest component, in this case, the E-O interface and digital computation. Alternatively, nonlinearities can also occur in the analog domain, leveraging the inherent nonlinear response of specific electronic components or analog circuits [122, 123, 124].
@@
Beyond implementing nonlinearity purely in the electronic domain, nonlinearity can also be introduced within the E-O/O-E conversion processes. Firstly, the activation function can be implemented by the nonlinear transmission properties of E-O modulators during the encoding. For instance, the nonlinear part of the cosine term in the MZM transfer function exhibits similarities to the sigmoid function. The methodology has been validated on the last two layers of an ONN for the MNIST classification task in Ref [104], and a similar approach involving the built-in nonlinearity of MRRs has been proposed by Gu, et al. [119]. The main advantage of introducing nonlinearity in encoding is it does not need extra O-E conversions, thus maintaining the inherent high bandwidth and speed of ONNs. However, since the nonlinearity is entirely contingent on the characteristics of E-O modulators, this dependency may pose challenges during the training phase. Inappropriate activation functions, particularly in deep networks, can result in problems such as gradient vanishing or explosion and low training efficiency. To achieve reconfigurable activation functions, Williamson et al. introduced a nonlinear unit that converts a small portion of the optical output into an electrical signal to modulate the original optical signal (Fig. 8.a) [125, 126]. This setup offers two approaches for the electrical processing part. The first approach converts the tapped signal into an electrical signal directly, allowing for moderate adjustments in nonlinearity through varying electrical biases. The second strategy employs reconfigurable lookup tables operated by a microcontroller, enabling the generation of arbitrary nonlinearities, aligning more closely with an all-electrical approach. Likewise, MRRs can be employed to realize this methodology. In Ref [127] and [128], activation functions are implemented through the modulation of a CW probe, utilizing the output from photodetectors (Fig. 8.b). While this architecture avoids the loss accumulation issue, it requires an additional laser source as the probe. In addition to introducing nonlinearity from modulation, O-E conversion processes, exemplified by the response of photodetectors, can also serve as sources of nonlinearity [129]. Nevertheless, these approaches are also constrained by the limitations of transfer characteristics.

### v2

@@ III-B1 Coherent vs incoherent ONN
Here, $\kappa$ and $t$ represent the field coupling factor and transmission factor, respectively. $\alpha_{rt}=exp\left(-\alpha\cdot 2\pi R\right)$ and $\phi_{rt}=\beta\cdot 2\pi R$ are the round-trip field attenuation factor and phase, with $\alpha$ and $\beta$ being the real and imaginary parts of the complex transmission coefficients [118]. The resonance condition is $\phi_{rt}=2m\pi$, i.e., $\lambda_{\text{res}}=2\pi n_{\text{eff}}R/m$, where $m$ is a positive integer. A potential issue in achieving full-range modulation (i.e., from 0 to 1) arises, as it can only be attained with symmetric, lossless coupling ($\alpha_{rt}=1$, $t_{1}=t_{2}$), which is unrealistic for real devices. An all-pass MRR can be regarded as a special case of an add-drop MRR, distinguished by the absence of a drop bus waveguide with $t_{2}=1$ and $\kappa_{2}=0$ for Eq. (16). While the phase at the through/drop port can also be derived from Eq. (16) and Eq. (17), the phase response and tuning range are highly sensitive to coupling conditions and round-trip loss. Hence, MRRs are primarily employed for amplitude modulation in incoherent systems. Another challenge comes from the experimental perspective: MRRs, particularly those with high quality factors, are sensitive to environmental factors such as temperature variations and vibrations. For a more in-depth theoretical analysis of MRRs, readers may refer to relevant literature and books [118, 119].
In tensor operations, MRRs can be modulated by tuning coupling coefficients or the round-trip phase [120]. Unlike MZIs, MRRs demonstrate significant wavelength selectivity. An MRR with a moderate quality factor can achieve a resonance peak with a 3 dB bandwidth of a few hundred picometers. The narrow bandwidth makes MRRs particularly effective as tunable filters in WDM-based PICs, where they are often arranged in series for the selective modulation of signals with different wavelengths. This application is exemplified by the concept of “weight banks” proposed by Tait et al. [121, 122], as shown in Fig. 6.d. In this architecture, photodetectors can spontaneously perform the sum of operands encoded on different wavelengths. Another incoherent architecture is the crossbar array, which implements MVMs using its programmable transfer matrix in the form of a tunable switch array [123, 124, 125]. Ohno, et al. demonstrated a 4$\times$4 add-drop MRR crossbar array for MVMs (Fig. 6.e) [123]. Specifically, each wavelength coming from the row bus waveguide can be weighted and subsequently directed into a specific column by the MRR array. The weighted elements can be collected and aggregated by a photodetector at the end of the column, thus carrying out MVM.
@@
Figure 6: Various implementations of photonic-electronic tensor cores. (a) Schematic of a 2$\times$2 MZI where the beam splitters can be implemented by multi-mode interferometer (MMIs) or directional couplers. An alternative configuration of MZI features a single input and output, with the splitter designed as a Y branch. (b) An MZI-based coherent photonics tensor core for ONNs [28]. The weight matrix can be represented by SVD in the form $\bm{U\Sigma V^{\dagger}}$. Due to chip size and complexity considerations, this work only implements $\bm{U}$ and $\bm{\Sigma}$ on a single pass through the circuit. (c) Schematic and power spectral responses of an add-drop MRR with power coupling coefficients ${\kappa_{1}}^{2}={\kappa_{2}}^{2}=0.2$ (assuming the coupling junctions are lossless, i.e., $\kappa^{2}+t^{2}=1$) and 5% round-trip power loss. (d) The weight bank architecture utilizes MRRs as tunable filters for parallel modulation of WDM signals [122]. The series-connected MRRs share a single bus waveguide as a through port, while the other waveguide can be shared or independent, enabling the selective addition or removal of wavelengths. (e) Schematic of 4$\times$4 MRR crossbar array demonstrated in Ref [123]. Input signals are first modulated via MZMs, and then distributed and weighted through the crossbar array, followed by a photodetector array that sums the weighted WDM elements. (f) The Schematic of an optical tensor core supports dynamic full-range general matrix multiplication based on WDM and coherent interference [126].
@@ III-B2 Static and dynamic weight encoding
The above-mentioned architectures topically map one operand of MVMs (typically a weight matrix W) onto hardware, executing multiplication through the transmission matrix of the PIC. Therefore, the operational speed of an ONN is, in theory, only determined by the modulation rate of input signals x. However, practical implementations face several limitations. Primarily, deploying MVMs with large dimensions is challenging due to the cost, complexity, loss, and other scalability issues of PICs. While multiplexing techniques can enhance the parallelism of ONNs, time-domain hardware reuse is necessary to fulfill the computational requirements of complex machine learning tasks. This indicates that the “weight-static” architecture constitutes a bottleneck for large networks because of the huge gap between ultra-fast optical computing and slow mapping/reprogramming. To fully unleash the potential of optical analog computing, the critical role of dynamic encoding for specific tasks should be recognized. For dynamically-operated ONNs, two essential requirements must be satisfied. First, the parameters in the network need to be high-speed reprogrammable for efficient hardware reuse. This requires that the modulators operate at gigahertz-level rates, as exemplified by those based on field-effect tuning mechanisms. Secondly, dynamic operation mandates that parameter mapping and output reading be conducted “directly” without additional signal preprocessing or data correction. For instance, to map a 12×12 matrix to an MZI-mesh framework requires $\sim$1.5 ms to perform SVD and phase decomposition on a CPU [126], which precludes dynamic operation due to the delay.
The importance of dynamic encoding is further highlighted by its compatibility with Transformer [9]. Transformer and the unique attention mechanisms have attracted significant research interest in recent years due to their exceptional performance in natural language processing (NLP), machine vision (MV), and large-scale language models (LLM). Unlike weight-static architectures, Transformer employs the multi-head self-attention (MHA) mechanism within its encoder and decoder blocks, necessitating matrix multiplication involving two dynamic, full-range operands. Zhu et al. presented a dynamically-operated PTC for the first photonic Transformer accelerator, leveraging coherent light interference and WDM [126]. As shown in Fig. 6.f, the elements within the input vectors x and y are encoded at distinct wavelengths, and then add a $-90^{\circ}$ phase shift to one vector. Through a directional coupler, two orthogonal signals can be recombined into the form (x $\pm$ y), facilitating the computation of the dot product via a balanced photodetector.
@@ III-B3 Hardware-efficient PTCs
For the aforementioned structures, the total number of optical components required to implement an $m\times n$ general matrix is similar, e.g., $m(m-1)/2+n(n-1)/2+\max(m,n)$ MZIs for an SVD-based MZI mesh, and $m\times n$ MRRs in microring-based ONNs. In addition to the high hardware costs and inherent control complexities, the required electrical components make up a large proportion of energy consumption, particularly for high-speed and high-resolution modulation. To address these challenges, researchers are exploring strategies to enhance hardware efficiency across multiple levels—from the device to the architecture—to improve the scalability of ONNs.
Firstly, optical devices or structures featuring compact topology have been proposed to reduce PIC footprint and improve hardware efficiency. Zhu et al. developed an ONN architecture consisting of two diffractive cells and MZIs exhibiting linear scaling in relation to input dimensions [127], as shown in Fig. 7.a. The star-shaped diffractive cell shows the capability of performing on-chip parallel Fourier transform and convolution operations. In a similar vein, an ONN leveraging the combination of WDM and MMIs has been developed and demonstrates an accuracy of 92.17% on the MNIST dataset. Beyond waveguide-based PIC, Wang et al. proposed a metasurface-based processing unit that provides ultra-high throughput for MVM [128] (Fig. 7.b). Within the subwavelength structure, each pair of slots acts as a weight element and connects to the following layers via in-plane diffraction and interference. For a specific neural network, the width and length of the slot need to be designed to map the corresponding weights.
The passive ONNs not only yield smaller footprints and enhanced hardware efficiency but also lower power consumption. However, a significant challenge associated with the passive task-specific PTCs is their fixed configuration, which can lead to degraded effectiveness when applied to varying tasks. To address this challenge, active multi-operand devices present another opportunity to break the fundamental limitation to achieve high-density tensor operation by squeezing MAC operation into a single device. Specifically, the execution of the length-k dot product between input vector $\bm{x}_{\text{in}}$ and weight vector $\bm{w}$ by k-operand devices can be represented as follows:
@@
where the function $f(\cdot)$ represents the E-O transmission function, while $g(w,x)$ is related to the encoding mechanism. Here, $\bm{w}$ can be encoded using programmable resistances (such as memristors or PCMs), tunable amplifiers/attenuators, or the effective modulation length in cases, while $\bm{x}_{\text{in}}$ can be encoded by electrical signals. The type of multi-operand devices can either be the modulators mentioned above or based on other passive optical components such as MMI [129]. Feng et al. first demonstrated this method experimentally on a 4-operand MZI, achieving a measured accuracy of 85.9% in SVHN recognition tasks with 4-bit control precision (Fig. 7.c). Gu et al. proposed a compact ONN architecture based on multi-operand MRRs (MOMRR) [130]. The MOMRR-based ONN supports weight encoding through two sets of rails, with full-range results carried out by balanced photodetectors [131]. Additionally, the architecture can be combined with the structured pruning strategy to further improve network scalability. Theoretically, multi-operand optical synapses could execute vector operations with nearly the same footprint as the single-operand counterparts. However, the intrinsic nonlinearity and possible crosstalk among operands could present challenges for training and calibration.
@@
Beyond improvements at the device level, another promising approach enhances hardware efficiency at the circuit level by software-hardware co-design approaches. Subspace neural networks, for instance, sacrifice a portion of matrix representability in exchange for fewer parameters instead of implementing universal linear operations or general matrix multiplication (GEMM). The effectiveness of this strategy can extend to ONNs by trading the universality of weight representation for higher hardware efficiency. An example is the butterfly-style PTC [89] (Fig. 7.d). By parameter pruning, this architecture could be implemented using significantly fewer MZIs with a scale factor of $O(nlog_{2}^{n})$ rather than the $O(n^{2})$ required by an MZI mesh for GEMM. This study experimentally showcases a measured accuracy of over 94% on the MNIST hand-written digits classification task, with up to 7$\times$ fewer active optical components, a 3.3× smaller footprint, and a 5.5$\times$ lower latency compared to conventional MZI mesh. A similar strategy, utilizing an MRR-based crossbar array to implement block-circulant matrices, has been demonstrated in Ref [125]. For subspace networks, the trade-off between hardware efficiency and matrix representability is an important consideration, and more details will be discussed in Section V-A.
@@
Beyond the one-shot broadcast architecture, an alternative method leverages time-domain integration with fewer optical components to perform MACs [132, 133]. Following this approach, Sludds et al. developed a delocalized optical computing system that showcases edge computing capabilities over a span of 86 kilometers [132]. As shown in Fig. 7.e, the modulated WDM signals are separated by a passive demultiplexer and then fed to a set of time-integrating receivers. On the client side, only one MZM, ADC, and DAC are used, which allows an ultra-low power assumption of femtojoules per MAC operation. The photocurrent $I(t)$ generated by the photodetector produces a voltage across the integrating capacitor $C$ by accumulating charge, thereby achieving the summation operation, which can be expressed as Eq. (19).
@@
This strategy sacrifices the speed benefit of optical computing to achieve a smaller chip footprint and reduced hardware complexity, which enables the use of milliwatt-class edge devices.
@@ III-C Implementations of Nonlinearities
Activation functions introduce the necessary nonlinearity into the network. For PIC-based photonic neurons, the implementations of nonlinearities fall into two major categories: the optical-electrical-optical (O-E-O) and the all-optical approach.
@@ III-C1 O-E-O nonlinearities
The activation function within O-E-O neurons can be realized most directly by routing the weighted sum through A/D conversion into digital processing units, such as CPUs, GPUs, or FPGAs. This process performs nonlinear operations digitally, and then converts the digital output back into analog signals that are subsequently fed into photonic neurons. The primary advantage of this approach is its extensive transfer ability of existing ANN architectures to the PIC platform, which also facilitates the implementation of arbitrary activation functions. However, the compulsory A/D and D/A conversion, along with digital processing, impose latency and power consumption, thereby undermining the efficiency advantage inherent in ONNs. As mentioned in Section II, the performance and speed of an optical computing system are constrained by its weakest link, which in this case is the E-O interface and the digital processing. Alternatively, nonlinearities can also occur in the analog domain by leveraging the inherent nonlinear responses of specific electronic components or analog circuits [134, 135].
@@
Beyond implementing nonlinearity purely in the electronic domain, nonlinearity can also be introduced within the E-O/O-E conversion processes, such as the nonlinear transmission properties of E-O modulators during encoding. For instance, the nonlinear part of the cosine term in the MZM transfer function exhibits similarities to the sigmoid function. This methodology has been validated on the last two layers of an ONN for the MNIST classification task in Ref [115], and a similar approach involving the built-in nonlinearity of MRRs has been proposed by Gu, et al. [131]. The main advantage of introducing nonlinearity in encoding is it does not need extra O-E conversions. However, since the nonlinearity is entirely contingent on the transmission characteristics of E-O modulators, this dependency may pose challenges during the training process. Inappropriate activation functions, particularly in deep networks, can result in gradient vanishing or explosion and low training efficiency. To achieve reconfigurable activation functions, Williamson et al. introduced a nonlinear unit that converts a small portion of the optical output into an electrical signal to modulate the original optical signal (Fig. 8.a) [136, 137]. This setup offers two approaches for the electrical part. The first approach converts the tapped signal into an electrical signal directly, allowing for moderate adjustments in nonlinearity through varying electrical biases. The second strategy utilizes reconfigurable lookup tables controlled by a microcontroller, which enables the generation of arbitrary nonlinearities and aligns more closely with an all-electrical approach. Likewise, MRRs can be employed to realize this methodology. In Ref [138] and [139], activation functions are implemented by modulating a CW probe with the output from photodetectors (Fig. 8.b). While this architecture avoids the loss accumulation issue, it requires an additional laser source as the probe. In addition to introducing nonlinearity from E-O modulation, O-E conversion processes, exemplified by the response of photodetectors, can also serve as the source of nonlinearity [140]. Nevertheless, these approaches are also constrained by limited reconfigurability.